# Spectraformer: A Unified Random Feature Framework for Transformer

### v2

@@ 3.3 Utility
There has been an unsystematic attempt at combining component functions and weight matrices. On one hand, most component functions have had theoretical and experimental results in combining with Base and ORF. On the other hand, most weight matrices have had results in combining with TrigRF or PosRF. This means that a lot of potential combinations have not been studied previously. Table 1 shows the wide research gap across 14 combinations which have never been studied in either the kernel or Transformer setting. Spectraformer highlights this research gap. Excluding TrigRF due to its instability and Base due to its underperformance, we conduct experiments on 18 combinations.
@@ 4.1 Experiment Details
The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to. Specifically, we evaluate the models on three LRA tasks: ListOps [30], Byte-Level Text Classification (using the IMDb review dataset) [31], and Byte-Level Document Retrieval (using the ACL Anthology Network dataset) [32]. All our models are experimented on Google Cloud Platform with n1-standard-4 (4 vCPUs, 15 GB RAM) and 1 NVIDIA T4 GPU. Our codebase is based on and includes the adapted or original implementation from [16], [14], [7], [8], [12], [13], [33] (Apache License, Version 2.0), [34] (Apache License, Version 2.0), [35], [36] (Apache License Version 2.0), [17].
Due to the significant number of models that are trained, we do not run the experiments on different seeds nor perform hyper-parameter tuning. The parameters are identical to [16] and chosen to limit the parameter count to account for the marked training time of multiple models. All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component. Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from [16]. Our hyperparameters are available in Table 5. The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
Table 2: Experimental results of Spectraformer variants on the LRA benchmark. We report accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)). $L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three. The best variant (PosRF-MM) with highest average accuracy, and overall low training time/memory is higlighted in red. Existing RF techniques are cited in the model column. Novel combinations are marked by $\checkmark$ on the Novel column.
@@
Novel Model L T R $\mu$ L T R $\mu$ L T R $\mu$
PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-SORF 22.98 63.31 65.52 50.60 1.29 2.81 2.83 2.31 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-QMC 37.50 60.41 80.71 59.54 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-SGQ 37.45 62.68 78.37 59.50 1.30 2.83 2.89 2.34 1.17 2.31 2.10 1.86
$\checkmark$ PosRF-MM 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-SORF 32.11 64.34 77.47 57.97 1.57 3.37 3.41 2.78 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-QMC 38.41 60.32 80.80 59.84 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-MM 38.71 60.39 80.45 59.85 1.61 3.46 3.51 2.86 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-SGQ 22.53 61.34 79.29 54.39 1.60 3.45 3.47 2.84 1.36 2.71 2.56 2.21
$\checkmark$ OPRF-FastFoodL 37.40 64.04 78.32 59.92 2.81 5.84 5.86 4.84 0.85 1.69 1.67 1.40
SADERF-ORF [13] 37.35 61.37 80.87 59.86 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-SORF 32.31 64.49 76.43 57.74 1.58 3.44 3.49 2.84 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-QMC 37.70 59.89 80.65 59.41 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-MM 38.00 60.80 80.48 59.76 1.61 3.52 3.58 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-SGQ 36.79 63.55 77.31 59.22 1.61 3.52 3.57 2.90 1.44 2.86 2.69 2.33
$\checkmark$ SADERF-FastFoodL 28.63 64.64 77.61 56.96 2.81 5.89 5.92 4.87 0.92 1.83 1.79 1.51
4.2 Comparison of Weight Matrix-Component Function Alternatives using Spectraformer
Our Spectraformer comparison results are available in Table 2. We now analyze the performance of component functions and weight matrices. We perform our analysis on both general trends of component functions across weight matrices and vice versa. We also make use of mean statistics (i.e., mean accuracy, mean training time, and mean peak memory consumption) which we give in brackets in the subsequent discussion.
Figure 3: Peak memory consumption (x-axis), accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval. Color indicates training time, the darker the faster, the brighter the slower (see legend). Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature. Values for non-RF Transformers are taken from [16].
Component functions. In terms of accuracy, SADERF (58.82%) typically performs better than OPRF (58.53%) and OPRF performs better than PosRF (57.50%) by small margin, agreeing with [12, 13]. However, SADERF (3.21 hours, 2.19 GB) and OPRF (3.17 hours, 2.24 GB) consistently take longer to train and have higher memory consumption than PosRF (2.71%, 1.76 GB) across weight matrices. This is likely due to OPRF and SADERF having additional computation of the L2 norm of the weight matrix which PosRF does not calculate.
Weight matrices. In terms of accuracy, SGQ (59.70%), ORF (59.64%), and QMC (59.59%) have the best performance. This is followed by MM (57.93%), FastFoodL (57.41%), and the worst being SORF (55.43%). In terms of training time, most weight matrices are rather uniform: SORF (2.64 hours) and MM, SGQ, QMC, ORF all at 2.69 hours. FastFoodL takes significant amount of time (4.78 hours). For peak memory consumption, FastFoodL is the most efficient at 1.73 GB whilst all other methods maintain a uniform value of 2.13 GB. FastFoodL reduces the space complexity at the cost of the time complexity. SORF reduces training time whilst compromising to a small extent its accuracy. This is interesting as it violates our initial assumptions that learners should perform better than approximators.
Best model. Overall, we find that PosRF-MM has the best performance, with the top accuracy (60.19%), top-3 training time (2.35 hours) and top-3 peak memory consumption (1.86 GB), outperforming the SOTA SADERF-ORF [13] (59.86%, 2.90 hours, and 2.33 GB) and OPRF-ORF [12] (59.25%, 2.86 hours, 2.21 GB) in the existing RF Transformer family. PosRF-MM achieves this performance due to the combination of both an efficient component function (PosRF) and performative weight matrix (MM). PosRF-MM has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption. Our result demonstrates the viability of Spectraformer as a unifying random feature framework of Transformer.
Table 3: Comparison of the best Spectraformer variant (PosRF-MM), highlighted in red, against the OT and other efficient Transformers (RF and non-RF) in the literature. We report three metrics: accuracy (on the test set), time (training time (hour)), and memory (peak memory consumption (GB)). $L$ stands for ListOps, $T$ for Text, $R$ for Retrieval, and $\mu$ for the average of the three. Best results for each section and task are boldened.
@@ 4.2 Comparison of Weight Matrix-Component Function Alternatives using Spectraformer
Model L T R $\mu$ L T R $\mu$ L T R $\mu$
Original Transformer [1] 38.37 61.95 80.69 60.34 2.24 4.30 8.33 4.96 5.37 10.37 10.77 8.84
Nyströmformer [17] 38.51 64.83 80.52 61.29 0.71 0.71 1.29 0.90 1.37 1.21 2.39 1.66
Linformer [35] 37.45 58.93 78.19 58.19 0.60 0.65 1.13 0.79 0.99 0.99 1.89 1.29
Informer [36] 32.53 62.64 77.57 57.58 1.19 1.60 2.91 1.90 4.85 5.12 5.77 5.25
Reformer [33] 37.68 62.93 78.99 59.87 0.85 0.94 1.73 1.17 1.61 1.61 2.98 2.07
BigBird [37] 39.25 63.86 80.28 61.13 1.88 2.00 3.81 2.56 2.71 2.83 4.97 3.50
Skyformer [16] 38.69 64.70 82.06 61.82 1.29 1.02 1.86 1.39 1.75 1.59 3.15 2.16
PosRF-ORF [8] 37.35 61.60 80.53 59.83 1.30 2.84 2.89 2.34 1.17 2.31 2.10 1.86
OPRF-ORF [12] 37.50 59.35 80.90 59.25 1.61 3.45 3.50 2.86 1.36 2.71 2.56 2.21
SADERF-ORF [13] 37.35 61.37 80.87 59.86 0.00 3.52 3.58 2.36 0.00 2.86 2.69 1.85
PosRF-FastFoodL [14] 24.95 64.76 76.37 55.36 2.69 5.59 5.61 4.63 0.78 1.56 1.53 1.29
Spectraformer (PosRF-MM) 38.05 61.85 80.67 60.19 1.31 2.84 2.89 2.35 1.17 2.31 2.10 1.86
4.3 Comparison with other efficient Transformers
When comparing with other non-random feature Transformers, we observe that PosRF-MM performs comparatively against the OT whilst reducing memory consumption and training time significantly (see Figure 3 and Table 3). On ListOps and Retrieval, PosRF-MM and other Spectraformer variants perform on par with other non-RF efficient Transformers. Results for non-RF efficient Transformers (above the horizontal divider) in Table 3 are taken from [16]. Spectraformer does not outperform more performant Nyström-based transformers which are not based on random features. Having said that, we are the first to have conducted a comprehensive analysis of random feature methods in the Transformer via the Spectraformer framework, comparing it against other past works and the OT. Spectraformer can be extended in the future to include a larger range of component function and weight matrix options. We discuss limitations, and broader impact of our research in Appendix A.1.
@@ 5 Conclusion
We present Spectraformer, a framework for approximating and learning the attention kernel in the Transformer. Our paper generalizes past works and presents empirical findings on different component function and weight matrix combinations. We experiment with 18 combinations, of which 14 are novel. Novel combinations outperform the previous SOTA in the family for multiple tasks. Our best model, PosRF-MM, has higher average accuracy (60.19) than the previous SOTA, SADERF-ORF (59.86), whilst having 23.4% faster training time, and 25.2% lower memory consumption. This shows the viability of Spectraformer as a unifying random feature framework of Transformer.

### v3

@@ 3.3 Utility
There has been an unsystematic attempt at combining component functions and weight matrices. On one hand, most component functions have had theoretical and experimental results in combining with Base and ORF. On the other hand, most weight matrices have had results in combining with TrigRF or PosRF. This means that a lot of potential combinations have not been studied previously. Table 2 shows the wide research gap across 14 combinations which have never been studied in either the kernel or Transformer setting. Spectraformer highlights this research gap. Excluding TrigRF due to its instability and Base due to its under-performance, we conduct experiments on 18 combinations.
@@
3.4 Adding New Code to Spectraformer
We now discuss how to add a new component function or weight matrix to the base code for further implementation. New component functions need to satisfy Definition 3.2 and new weight matrices need to satisfy Definition 3.1.
3.4.1 Component Function
  1. 1.
    Add the new component function f to src/models/component_functions.py, the arguments should include data (the input), and other optional parameters.
  2. 2.
    Import f and add a new entry to FastAttention.comp_functions[f_name] = f in src/models/attention_performer.py (line 176)
3.4.2 Weight Matrix
  1. 1.
    Add the new weight matrix w to src/models/weight_matrix_approx.py or src/models/weight_matrix_learner.py, the arguments should include nb_rows (number of rows), nb_cols (number of columns) and device.
  2. 2.
    Import w and add the if clause and $w$ function call in src/models/attention_performer.py (line 208)
@@ 4.1 Experiment Details
Table 3: Experimental results of Spectraformer variants on the LRA benchmark, over five seeds. We report mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)). $L$: ListOps, $T$: Text, $R$: Retrieval, and $\mu$ for the average of the three. Entries are ranked by $\mu$ accuracy. The previous RF SOTAs (SADERF-ORF (Likhosherstov et al., 2023) and OPRF-ORF (Likhosherstov et al., 2022)) are in bold. We note that via Spectraformer, we find various variants outperforming these previous RF SOTAs in terms of accuracy, with different combinations offering different trade-off between accuracy, time, and memory. Entries are sorted by mean accuracy.
@@
L T R $\mu$ L T R $\mu$ L T R $\mu$
OPRF-FastFoodL 37.55 (0.48) 64.41 (0.62) 77.70 (0.33) 59.89 1.07 2.07 2.12 1.75 0.86 1.72 1.68 1.42
OPRF-MM 38.08 (0.53) 60.40 (0.85) 81.09 (0.18) 59.86 0.68 1.26 1.24 1.06 1.36 2.71 2.56 2.21
PosRF-MM 37.06 (0.37) 61.87 (1.79) 80.58 (0.53) 59.84 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
OPRF-ORF 38.34 (0.22) 60.16 (0.79) 80.88 (0.17) 59.80 0.68 1.26 1.25 1.06 1.36 2.71 2.56 2.21
SADERF-QMC 37.37 (0.38) 61.14 (1.48) 80.84 (0.11) 59.78 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
PosRF-QMC 37.11 (0.09) 61.69 (0.96) 80.55 (0.13) 59.78 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
SADERF-MM 37.10 (0.22) 60.68 (1.88) 81.13 (0.17) 59.64 0.68 1.25 1.29 1.07 1.44 2.86 2.69 2.33
SADERF-ORF 37.10 (0.19) 60.39 (2.08) 81.05 (0.22) 59.51 0.68 1.25 1.28 1.07 1.44 2.86 2.69 2.33
OPRF-QMC 37.69 (0.62) 59.94 (0.59) 80.38 (0.49) 59.34 0.68 1.26 1.26 1.07 1.36 2.71 2.56 2.21
SADERF-SGQ 37.11 (0.21) 62.46 (0.54) 78.38 (0.25) 59.32 0.68 1.25 1.27 1.07 1.44 2.86 2.69 2.33
SADERF-FastFoodL 36.02 (1.38) 64.63 (0.18) 76.99 (0.61) 59.21 1.07 2.08 2.16 1.77 0.92 1.84 1.80 1.52
OPRF-SGQ 37.10 (0.23) 61.25 (0.54) 78.69 (0.54) 59.01 0.67 1.26 1.25 1.06 1.36 2.71 2.56 2.21
PosRF-ORF 34.35 (5.96) 60.30 (0.97) 80.45 (0.22) 58.37 0.56 1.05 1.06 0.89 1.17 2.31 2.10 1.86
PosRF-FastFoodL 33.46 (3.70) 64.65 (0.36) 76.95 (0.48) 58.35 1.02 1.98 2.03 1.68 0.79 1.57 1.53 1.30
SADERF-SORF 33.30 (0.98) 64.70 (0.36) 74.71 (1.90) 57.57 0.68 1.24 1.28 1.07 1.44 2.86 2.69 2.33
PosRF-SGQ 28.64 (7.54) 62.38 (0.53) 78.28 (0.20) 56.43 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
OPRF-SORF 27.91 (3.26) 64.76 (0.66) 75.92 (1.74) 56.20 0.67 1.26 1.24 1.06 1.36 2.71 2.56 2.21
PosRF-SORF 21.27 (6.65) 62.99 (0.40) 67.10 (1.11) 50.45 0.56 1.05 1.05 0.89 1.17 2.31 2.10 1.86
The LRA covers tasks of different sequence lengths, difficulty, and objective and it is designed to evaluate efficient Transformers, of which Spectraformer belongs to. Specifically, we evaluate the models on three textual LRA tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013). We discuss the task characteristics in greater details in Section 4.2. All our models are experimented on NVIDIA V100 and 12 CPU with 20GB of memory. Our codebase is based on and includes the adapted or original implementation from Chen et al. (2021); Chowdhury et al. (2022); Liu et al. (2022); Choromanski et al. (2021); Likhosherstov et al. (2022; 2023); Wang et al. (2020); Xiong et al. (2021), Kitaev et al. (2020); Thomas et al. (2018); Zhou et al. (2021) (Apache License, Version 2.0).
We run our experiments on 5 different seeds due to computational limitation. The parameters are identical to Chen et al. (2021) and chosen to limit the parameter count to account for the marked training time of multiple models. All the parameters for Spectraformer variants are kept identical for fair comparison, with the only difference being their random feature component. Our code is implemented in Python 3.12 and Pytorch, we use the Transformer base from Chen et al. (2021). Our hyperparameters are available in Table 8. The model name convention is [component function]-[weight matrix] (e.g., OPRF-FastFoodL).
Our result in Table 3 shows that the top mean accuracy models, ranked in descending order, are: OPRF-FastFoodL, OPRF-MM, PosRF-MM, OPRF-ORF, SADERF-QMC. Our mean accuracy result indicates that novel Spectraformer combinations have outperformed the highlighted previous SOTA, i.e., OPRF-ORF and SADERF-ORF. However, we are aware that mean accuracy is insufficient in offering a comprehensive picture. Therefore, we first proceed with analyzing the statistics of the tasks before analyzing the significance of these results.
4.2 Task Characteristics
We first analyze the characteristics of the three textual classification LRA (Tay et al., 2021) tasks: ListOps (Nangia & Bowman, 2018), Text (Byte-Level Text Classification, using the IMDb review dataset) (Howard & Ruder, 2018), and Retrieval (Byte-Level Document Retrieval, using the ACL Anthology Network dataset) (Radev et al., 2013). We take into consideration three main characteristics:
  • Number of output classes. ListOps is a task which inputs a sequence containing brackets and simple mathematical operators, with the output being a integer between 1 and 10, hence it has 10 classes. Text is a sentiment classification task on an IMDb review dataset with the class being either negative or positive, hence it has 2 classes. Retrieval inputs two sequences and the output is either False (meaning the sequences are not of the same document) or True (meaning the sequences are of the same document).
  • Maximum sequence size. ListOps is 2K; Text is 4K; whilst Retrieval requires the processing of two sequences each of 4K length making the total maximum sequence size being 8K.
  • Required attention span. This is the ‘mean distance between the query token and the attended tokens, scaled by attention weights’. The average values for required attention span calculated by Tay et al. (2021) is: ListOps being 0.8, Text being 0.3, and Retrieval being 1.3.
Table 4: Mean Pearson correlation between model performance statistics and each task characteristic across weight matrices and component functions. Model performance statistics include ‘acc’ (accuracy), ‘time’ (training time), ‘mem’ (peak memory consumption). Task characteristics are ‘size’ (maximum sequence size), ‘span’ (required attention span), and ‘classes’ (number of output classes). Correlation is calculated for each weight matrix or component function and the average value over weight matrix or component function respectively displayed. Left is weight matrix $W$, right is component function $f$.
$W$ acc time mem
size 0.93 0.74 0.64
span 0.37 0.02 -0.10
classes -0.91 -0.96 -0.94
$f$ acc time mem
size 0.93 0.56 0.60
span 0.36 0.02 -0.10
classes -0.91 -0.72 -0.89
We then calculate Pearson correlation of accuracy, training time, and memory consumption against these characteristics. Our result is shown in Table 4. The table indicates that: (1) contrary to intuition, longer sequence size gives higher accuracy but this is not because of increasing required attention span and they somewhat increase training time and peak memory consumption; (2) increasing the number of classes decreases accuracy, training time, and peak memory consumption.
4.3 Analysis of Combinations
Table 5: Mean statistics across weight matrices for each task. Task is denoted on the top left cell of each table. Mean statistics include ‘acc’ (accuracy (%)), ‘time’ (training time (hour)), and ‘mem’ (peak memory consumption (GB)).
ListOps acc time mem
SORF 27.49 0.64 1.32
SGQ 34.28 0.64 1.32
FastFoodL 35.68 1.05 0.86
ORF 36.59 0.64 1.32
QMC 37.39 0.64 1.32
MM 37.41 0.64 1.32
Text acc time mem
ORF 60.28 1.19 2.63
QMC 60.92 1.19 2.63
MM 60.98 1.19 2.63
SGQ 62.03 1.19 2.63
SORF 64.15 1.18 2.63
FastFoodL 64.56 2.04 1.71
Retrieval acc time mem
SORF 72.58 1.19 2.45
FastFoodL 77.22 2.10 1.67
SGQ 78.45 1.19 2.45
QMC 80.59 1.20 2.45
ORF 80.79 1.20 2.45
MM 80.93 1.20 2.45
We now explore the suitability of weight matrices and component functions depending on task characteristics. Table 5 shows the mean statistics across weight matrices for each task. We notice that weight matrices form two different groups which align with the required attention span differences. ListOps and Retrieval are tasks with high required attention span, ListOps being 0.8 and Retrieval being 1.3. In both these tasks, we observe that ORF, QMC, and MM perform comparatively against each other and a lot better against SORF, SGQ, and FastFoodL. However, the Text task, which has a much lower required attention span of 0.3, has much higher performance from the SORF-SGQ-FastFoodL group, whilst ORF, QMC, and MM perform much poorly on this task. Based on this observation, we can conclude that ORF, QMC, and MM are suited for high required attention span; whilst SORF, SGQ, and FastFoodL are suited for low required attention span. Analyzing mean statistics per task across component functions do not show any task-specific trend (refer to Table 9 for more information).
Table 6: Summary statistics of component functions $f$ (top) and weight matrices $W$ (bottom) in mean accuracy (on the test set), mean time (training time (hour)), and mean memory (peak memory consumption (GB)). $f$ and $W$ entries are sorted by mean accuracy respectively.
@@ 4.3 Analysis of Combinations
$f$ SADERF 59.17 1.19 2.19
$f$ OPRF 59.02 1.18 2.08
$f$ PosRF 57.20 1.02 1.77
$W$ MM 59.78 1.01 2.13
$W$ QMC 59.63 1.01 2.13
$W$ ORF 59.22 1.01 2.13
$W$ FastFoodL 59.15 1.73 1.41
$W$ SGQ 58.25 1.01 2.13
$W$ SORF 54.74 1.00 2.13
In analyzing efficiency, we follow the definition of efficient Transformers in Tay et al. (2022) in terms of memory and computation, the latter interpreted as computational cost. The overall trend of accuracy, time, and memory of component functions is that they increase linearly from PosRF, OPRF, to SADERF (see Table 6). Hence, when limited computational resource is a concern, PosRF is a better option. Training time and peak memory consumption relies more heavily on component functions than weight matrices, with the exception of FastFoodL. FastFoodL takes significantly less memory consumption at the expense of training time. The accuracy trend for weight matrices is as described above. Table 6 also shows that whilst FastFoodL generally does not work as well due to its low required attention span, its high performance in the task of Text which is low in required attention span and the anomalously performant OPRF-FastFoodL leads to a data skew in the final ranking in Table 3.
Comparing between combinations, we find that for ListOps, OPRF-ORF is the most accurate (38.34%), followed by OPRF-MM (38.08%) and OPRF-QMC (37.69%). OPRF-ORF, being the previous SOTA, is undoubtedly the best choice here since other models maintain the same training time and peak memory consumption with lower accuracy. For Text, the most accurate model is OPRF-SORF (64.76%), followed by SADERF-SORF (64.70%) and PosRF-FastFoodL (64.65%). These all outperform the previous SOTAs (OPRF-ORF and SADERF-ORF) by $4\%+$ whilst maintaining a diversity of training time and peak memory consumption options. Both OPRF-SORF and SADERF-SORF maintain the same level of training time and peak memory consumption as the SOTAs, PosRF-FastFoodL offers much less peak memory consumption at the expense of longer training time. For Retrieval, the most accurate model is SADERF-MM (81.13%), followed by OPRF-MM (81.09%) and SADERF-ORF (81.05%). Whilst these models only perform slightly better than the SOTA SADERF-ORF, OPRF-MM offers less peak memory consumption compared to SADERF-ORF.
@@ 5 Conclusion
We present Spectraformer, a framework for approximating and learning the attention kernel in the Transformer. Our paper generalizes past works and presents empirical findings on different component function and weight matrix combinations. We experiment with 18 combinations, of which 14 are novel. Novel combinations perform competitively against, if not more performant than the previous SOTA in the family for LRA textual tasks. We show empirically that the choice of performant weight matrix depends on the required attention span of the task. Our results and the extensibility of our work shows the viability of Spectraformer as a unifying random feature framework of Transformer.